OpenStudy (anonymous):

The radioactive substance strontium-90 has a half-life of 28 years. In other words, it takes 28 years for half of a given quantity of strontium-90 to decay to a non-radioactive substance. The amount of radioactive strontium-90 still present after t years is modeled by the expression grams. Evaluate the expression for t = 0, t = 28 and t = 56. What does each value of the expression 8(2^-t/28) represent?

OpenStudy (anonymous):

Do you know the formula for radioactive decay?

OpenStudy (anonymous):

no @geoffb

OpenStudy (anonymous):

8(2^-t/28) @geoffb ?

OpenStudy (anonymous):

Right. The question is asking what each value in that equation represents.

OpenStudy (anonymous):

so how do i answer that?

OpenStudy (anonymous):

Start by pulling out each value. 8 is a value. t is a value. 28 is a value. Et cetera.

OpenStudy (anonymous):

What do each of those represent?

OpenStudy (anonymous):

Uh, . . . that's not what's going on here, @geoffb

OpenStudy (anonymous):

It's the last sentence in the question. I figured that would be a better place to start, since he didn't know the formula. From there, he can plug in values of t and determine Q.

OpenStudy (anonymous):

" Evaluate the expression, 8(2^-t/28), for t = 0, t = 28 and t = 56." The formula was given in the problem statement.

OpenStudy (anonymous):

"What does each value ... represent?" means to interpret the output values after evaluating.

OpenStudy (anonymous):

Ahhh, yes. That makes more sense

OpenStudy (anonymous):

so i take this equation 8(2^-t/28) and change it to 8(2^-0/28), 8(2^-28/28), and 8(2^-56/28)??

OpenStudy (anonymous):

I'd recommend a T-table like this: 

OpenStudy (anonymous):

Yes, @chicagochica5 exactly!

OpenStudy (anonymous):

@CliffSedge and do i put them all negative since it was -t, or do i change it to positive?

OpenStudy (anonymous):

Yes, the - makes those values negative before you divide them by 28.

OpenStudy (anonymous):

@CliffSedge, for the past hour ive been working the problem and i dont know how to get the answers :( please help

OpenStudy (anonymous):

\[A(t)=8(2^{-t/28})\] First find A(0), A(28), and A(56) \[A(0)=8(2^{-(0)/28})\] \[A(28)=8(2^{-(28)/28})\] \[A(56)=8(2^{-(56)/28})\]

OpenStudy (anonymous):

@chicagochica5 were you able to find those 3 amounts yet?

OpenStudy (anonymous):

@CliffSedge no i cant solve them

OpenStudy (anonymous):

Simplify the exponents: \[A(0)=8(2^0)\]\[A(28)=8(2^{−1})\]\[A(56)=8(2^{−2})\] Does that help?

OpenStudy (anonymous):

the first one, No solution? the next one \[A = \frac{ 2 }{ 7 }\] ?

OpenStudy (anonymous):

and the last one \[A= \frac{ 1 }{ 28 }\] ? @CliffSedge

OpenStudy (anonymous):

Can you try \[2^{1}\] and \[2^{2}\]

OpenStudy (anonymous):

where

OpenStudy (anonymous):

A negative exponent is the reciprocal of the positive exponent. So, if you solve 2^1 and 2^2, take the reciprocal (one over the result) to get 2^-1 and 2^-2.

OpenStudy (anonymous):

Anything (except zero) to the 0th power = 1. Negative exponents mean to divide by that power.

OpenStudy (anonymous):

I'll just let my buddy, geoffb take it from here. ;-)

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

Okay sorry. I'll try again.

OpenStudy (anonymous):

its not you, im not good at math

OpenStudy (anonymous):

Do you know how to solve a number "to the power of" something, when it is positive? For example, do you know how to do 2 to the power of 2? \[2^{2}\]

OpenStudy (anonymous):

is 4?

OpenStudy (anonymous):

Yes, it is, but how did you figure that out? What exactly did you do?

OpenStudy (anonymous):

2 x 2

OpenStudy (anonymous):

Exactly. So what would be: \[2^{3}\]?

OpenStudy (anonymous):

And I don't need the answer—what would you do to *get* the answer?

OpenStudy (anonymous):

2 x 2 x 2

OpenStudy (anonymous):

Okay, good! You have a good grasp of positive exponents. Now...

OpenStudy (anonymous):

Do you know what a reciprocal is?

OpenStudy (anonymous):

a fraction flipped around

OpenStudy (anonymous):

Yes, exactly! Good. Now, do you know that every number can be represented as a fraction? For example, what would be the number 6 as a fraction?

OpenStudy (anonymous):

6/1?

OpenStudy (anonymous):

Yes, great! Which means you can always take the reciprocal by putting 1 over the number, because all you're doing is "flipping" the fraction. Does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Awesome. Now, as Cliff was about to say, a negative exponent simply means you take the reciprocal. So, the following are exactly the same. \[2^{-3} = \frac{1}{2^{3}}\] Still with me?

OpenStudy (anonymous):

wheres that equation from? how did you get that

OpenStudy (anonymous):

It's not from anywhere. I just picked it out of a hat. I could have said: \[x^{-2} = \frac{1}{x^{2}}\] or \[6^{-8} = \frac{1}{6^{8}}\]

OpenStudy (anonymous):

The point I'm trying to get across is that a negative exponent simply means you "flip the fraction" and find the reciprocal.

OpenStudy (anonymous):

oh ok, so 2^-3 = 1 over 2^3?

OpenStudy (anonymous):

Yes! Awesome.

OpenStudy (anonymous):

Now, going back to what Cliff wrote before, can you figure out these?: \[A(0)=8(2^{0})\] \[A(28)=8(2^{−1})\] \[A(56)=8(2^{−2})\]

OpenStudy (anonymous):

no i dont know how to solve

OpenStudy (anonymous):

Okay. You seemed to a minute ago. Be confident in yourself! :)

OpenStudy (anonymous):

If you had to take a wild guess, what would it be?

OpenStudy (anonymous):

for the first one, i had No solution

OpenStudy (anonymous):

The first one is actually, as Cliff said, any number (except 0) to the power of 0 is 1.

OpenStudy (anonymous):

thats the answer?

OpenStudy (anonymous):

Let's cut to the chase, so we can get to interpreting the values. 2^0 = 1 2^-1 = 1/2 2^-2 = 1/4

OpenStudy (anonymous):

ok and those numbers mean

OpenStudy (anonymous):

Multiplying those by the 8 grams you started with yields how much of the strontium is left after that amount of time.

OpenStudy (anonymous):



OpenStudy (anonymous):

multiply 0 x 8?

OpenStudy (anonymous):

No, t is the number of years elapsed. When t = 0, you have 8 grams.